OpenStudy (anonymous):

y=-1/2x+6 how do you know which one is the y intercept and the x intercept

OpenStudy (anonymous):

Put y=0 So, 0=-1/2x+6 or 1/2x=6 or 1=12x or x=1/12 Hence, x=1/2 is the x intercept. Similarly, put x=0 to get y intercept which is 6

OpenStudy (anonymous):

ok so if im looking for the y and x intercept of y=3x+5 the 5 is the y intercept and the x intercept is 3

OpenStudy (anonymous):

for y let x be zero y=3(0)+5 y=5 For x let y be zero 0=3x+5 -5=3x x=-5/3